OpenStudy (anonymous):

(cos^3x-sin^3x)/(cosx-sinx) = 1+sinxcosx Proving Trig Identity

OpenStudy (anonymous):

\[(\cos^3x-\sin^3x)/(cosx-sinx) = 1+sinxcosx\]

OpenStudy (precal):

multiply (cosx+sinx)/(cosx+sinx) to both sides

OpenStudy (anonymous):

what will that do? :/

OpenStudy (anonymous):

factoring the difference of two cubes might be a start

OpenStudy (anonymous):

how would i factor them?

OpenStudy (turingtest):

\[(a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

\[a^3-b^3=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

oh like that. okay I'll try that thanks.

OpenStudy (anonymous):

then cancel